Amore:

A car with a mass of 900 kg is moving with a speed of 60 kph. Calculate the kinetic energy of the car in J (3 points). A) 125000 J B) 7500 J C) 27000 J D) 1620000 J Please provide solution

Vocaloid:

kinetic energy = (1/2)mv^2 **however** since our energy unit is in joules, we have to convert kph to m/s. \[\frac{ 60~km }{ 1~hr }\frac{ 1~m }{ 1000km }\frac{ 1~hr }{ 3600s }\] notice how km and hours cancel out and you're left with m/s once you have the speed in m/s, plug into the KE formula